# Observational appearance and additional photon rings of the asymmetric thin-shell wormhole in the Kalb-Ramond background

### v1

@@ Abstract
Kalb-Ramond, asymmetric thin-shell wormhole, photon rings, Lorentz-violating
@@ I Introduction
The study of black hole shadows has always been a hot topic in the field of physics Perlick:2021aok ; Zhang:2024lsf ; Hou:2022gge ; Wang:2025ihg ; He:2024amh . Synge was the first to calculate and determine the boundary of the black hole shadow , while Luminet, pioneered the study of the angular radius equation for the photon capture region in a Schwarzschild black hole Synge:1966okc ; Luminet:1979nyg . In recent years, this field has seen remarkable developments within the framework of modified gravity theories. On the observational front, Heino Falcke and his colleagues were the first to propose the concept of the black hole shadow, and their work directly contributed to the development and advancement of EHT Falcke:1999pj . Although the EHT’s observations are largely consistent with the black hole shadow predictions of General Relativity, they do not rule out the existence of certain Ultra-Compact Objects (UCOs) Bambi:2008jg ; Banerjee:2019nnj ; Abdikamalov:2019ztb . In fact, objects such as wormholes and boson stars could potentially cast shadows very similar to those of black holes Tsukamoto:2021fpp ; Wielgus:2020uqz ; Wang:2023nwd ; Rosa:2023qcv ; Rosa:2022tfv ; Cunha:2016bjh . Consequently, finding methods to effectively distinguish UCOs from black holes has become a crucial research direction.
@@
Interestingly, Matt Visser proposed a novel construction known as the Thin-Shell Wormhole, which connects two distinct space-times via a throat, allowing for potential traversal Visser:1989kh ; Visser:1989kg . Building upon this theoretical framework, Wang investigates the optical appearance of an Asymmetric Thin-Shell Wormhole (ATSW) Wang:2020emr , and reveale that its shadow is consistently smaller than that produced by a black hole-a significant result that provides favorable support for the direct observation of wormholes. Subsequent research has primarily focused on the double shadow phenomenon generated by ATSWs. The presence of this double shadow may offer an alternative method for indirectly observing ATSW Guo:2022iiy ; Chen:2022tog . Observational results of these dual shadows could not only test predictions of General Relativity but also enhance our understanding of gravitational effects and space-time curvature Guerrero:2021pxt ; Guo:2022nto ; Wei:2023fkn . Furthermore, Peng et al. concentrated on the observational signatures of accretion disks and the additional photon rings around a Schwarzschild ATSW Peng:2021osd . These features can likewise serve as a method to distinguish ATSWs from black holes. Beyond this, other researchers have studied the observational appearance of wormholes in Hayward space-time and the appearance of a freely falling star within an ATSW Jing:2024thh ; Chen:2024wtw ; He:2024yeg ; Luo:2023wru . Inspired by these studies, this paper will investigate the observational characteristics of an ATSW surrounded by an accretion disk within the context of Kalb-Ramond(KR) gravity.
@@ II Null Geodesics and Effective Potential of the ATSW in the Kalb-Ramond Field
In the theoretical framework, $R$ denotes the Ricci scalar, $\eta$ and $\xi_{1}$, $\xi_{2}$ are coupling constants representing the non-minimal coupling terms between gravity and the KR field, while $\mathcal{L}_{\mathrm{N}}$ refers to the Lagrangian of the electromagnetic field. In our previous work, we investigated the influence of the charge $Q$ and the Lorentz-violating parameter $l$ on black hole shadows Tan:2025pya . Specifically, the study focused on a KR black hole surrounded by a thin accretion disk. It was found that an increase in the charge $Q$ and the Lorentz-violating parameter $l$ leads to a reduction in the size of the black hole shadow, as well as in the direct image, lensing ring, and photon ring. The present research shifts focus to the optical characteristics of an Asymmetric Thin-Shell Wormhole within the KR framework. This investigation is motivated by two key aspects: First, previous studies on Bardeen-type ATSWs have shown that the magnetic charge $g$ significantly affects their optical appearance He:2024yeg . This intriguing result prompts us to examine whether the charge $Q$ and the Lorentz-violating parameter $l$ similarly induce changes in the optical appearance of the KR field ATSW, analogous to the case of the KR black hole. Second, we aim to determine whether shadow-related characteristics can provide a viable method for distinguishing an ATSW from a black hole.
@@
That is, the metric function can be expressed as
@@
with $p_{i}^{r_{i}}=\mathrm{d}x^{r_{i}}/\mathrm{d}\lambda$ represents the four-momentum of the photon in the space-time, and $\lambda$ denotes the affine parameter. This affine parameter can be redefined as $\lambda=\lambda/E$, and $b$ is the impact parameter, which can be expressed as $b=|L|/E$. From the above relations, we derive the following

### v2

# Asymmetric thin-shell wormholes in the Kalb-Ramond background: Observational characteristics and extra photon rings

@@ Abstract
Kalb-Ramond, asymmetric thin-shell wormhole, observational characteristics, Lorentz-violation
@@ I Introduction
The study of black hole shadows has always been a hot topic in the field of physics Perlick:2021aok ; Zhang:2024lsf ; Zeng:2025olq ; Hou:2022gge ; Wang:2025ihg ; Zhang:2025vyx ; He:2024amh ; Fan:2024rsa . Synge was the first to calculate and determine the boundary of the black hole shadow , while Luminet, pioneered the study of the angular radius equation for the photon capture region in a Schwarzschild black hole Synge:1966okc ; Luminet:1979nyg . In recent years, this field has seen remarkable developments within the framework of modified gravity theories. On the observational front, Heino Falcke and his colleagues were the first to propose the concept of the black hole shadow, and their work directly contributed to the development and advancement of EHT Falcke:1999pj ; Zeng:2025kqw . Although the EHT’s observations are largely consistent with the black hole shadow predictions of General Relativity, they do not rule out the existence of certain Ultra-Compact Objects (UCOs) Bambi:2008jg ; Banerjee:2019nnj ; Abdikamalov:2019ztb . In fact, objects such as wormholes and boson stars could potentially cast shadows very similar to those of black holes Tsukamoto:2021fpp ; Wielgus:2020uqz ; Wang:2023nwd ; Rosa:2023qcv ; Rosa:2022tfv ; Cunha:2016bjh . Consequently, finding methods to effectively distinguish UCOs from black holes has become a crucial research direction.
@@
Interestingly, Matt Visser proposed a novel construction known as the Thin-Shell Wormhole, which connects two distinct space-times via a throat, allowing for potential traversal Visser:1989kh ; Visser:1989kg . Building upon this theoretical framework, Wang investigates the optical appearance of an Asymmetric Thin-Shell Wormhole (ATSW) Wang:2020emr , and reveale that its shadow is consistently smaller than that produced by a black hole-a significant result that provides favorable support for the direct observation of wormholes. Subsequent research has primarily focused on the double shadow phenomenon generated by ATSWs. The presence of this double shadow may offer an alternative method for indirectly observing ATSW Guo:2022iiy ; Chen:2022tog . Observational results of these dual shadows could not only test predictions of General Relativity but also enhance our understanding of gravitational effects and space-time curvature Guerrero:2021pxt ; Wang:2024uda ; Guo:2022nto ; Huang:2024wpj ; Wei:2023fkn . Furthermore, Peng et al. concentrated on the observational signatures of accretion disks and the additional photon rings around a Schwarzschild ATSW Peng:2021osd . These features can likewise serve as a method to distinguish ATSWs from black holes. Beyond this, other researchers have studied the observational appearance of wormholes in Hayward space-time and the appearance of a freely falling star within an ATSW Jing:2024thh ; Chen:2024wtw ; He:2024yeg ; Luo:2023wru . Inspired by these studies, this paper will investigate the observational characteristics of an ATSW surrounded by an accretion disk within the context of Kalb-Ramond(KR) gravity.
@@ II Null Geodesics and Effective Potential of the ATSW in the Kalb-Ramond Field
In the theoretical framework, $R$ denotes the Ricci scalar, $\eta$ and $\xi_{1}$, $\xi_{2}$ are coupling constants representing the non-minimal coupling terms between gravity and the KR field, while $\mathcal{L}_{\mathrm{N}}$ refers to the Lagrangian of the electromagnetic field. In our previous work, we investigated the influence of the charge $Q$ and the Lorentz-violating parameter $l$ on black hole shadows Tan:2025pya . Specifically, the study focused on a KR black hole surrounded by a thin accretion disk. It was found that an increase in the charge $Q$ and the Lorentz-violating parameter $l$ leads to a reduction in the size of the black hole shadow, as well as in the direct image, lensing ring, and photon ring. The present research shifts focus to the optical characteristics of an Asymmetric Thin-Shell Wormhole within the KR framework. This investigation is motivated by two key aspects: First, previous studies on Bardeen-type ATSWs have shown that the magnetic charge $g$ significantly affects their optical appearance He:2024yeg . This intriguing result prompts us to examine whether the charge $Q$ and the Lorentz-violating parameter $l$ similarly induce changes in the optical appearance of the KR field ATSW, analogous to the case of the KR black hole Zeng:2024ptv . Second, we aim to determine whether shadow-related characteristics can provide a viable method for distinguishing an ATSW from a black hole.
@@
The metric function can be expressed as
@@
where $p_{i}^{r_{i}}=\mathrm{d}x^{r_{i}}/\mathrm{d}\lambda$ represents the four-momentum of the photon in the space-time, and $\lambda$ denotes the affine parameter. This affine parameter can be redefined as $\lambda=\lambda/E$, and $b$ is the impact parameter, which can be expressed as $b=|L|/E$. From the above relations, we derive the following